OpenStudy (anonymous):

M is the average of 10 positive integers, from 1 to 100, inclusive. Please state what must be true of M. A. M is not an integer B. M is even C. M is odd D. M >= 6 E. M <= 96

OpenStudy (anonymous):

A) it could certainly be an integer right? \[\frac{10+20+30+40+50+60+70+80+90+100}{10}\]

OpenStudy (anonymous):

Yes. It could be an integer, so A is certainly not the correct answer. @satellite73

OpenStudy (loser66):

\(\dfrac{1+3+5+7+9+11+13+15+17+19}{10}=10\)

OpenStudy (loser66):

so c is not true, the same with b

OpenStudy (loser66):

But we need more given condition: does the number repeat?

OpenStudy (anonymous):

The answer is either D or E. You can figure out that the maximum value is 95.5, and the minimum value is 5.5 However, the answers are as provided. The numbers do not repeat.

OpenStudy (anonymous):

When the minimum value is 5.5, and the maximum value is 95.5, and they want you to state that M is either greater than or equal to 6, or M is less than or equal to 96, which do you choose as the correct answer?

OpenStudy (loser66):

Since you have the minimum is 5.5 , so the so-called "Must be greater or equal 6" sounds not good to me.

OpenStudy (anonymous):

I was thinking that the answer would be E. Still pisses me off though, since m CAN'T be equal to 96.

OpenStudy (loser66):

\(\leq\) means lesser than OR equal, not AND

OpenStudy (loser66):

so that the result can be lesser OR equal 96, if it can't be 96, it can be lesser. It's true still